OpenStudy (anonymous):

If a cylinder holds a volume of 100 cm3, and is 30 cm tall, which is the area of the bottom of the cylinder?

OpenStudy (anonymous):

Volume of cylinder, V\[=\pi r ^{2} h\] Area of circle\[=\pi r ^{2}\] =V/h =100/30 \[\approx 3.33 cm\]

OpenStudy (anonymous):

Next time, please post these types of questions to the Mathematic group. They can be misleading sometimes. ;)